OpenStudy (anonymous):

WRITE THE RATIO AS A FRACTION IN LOWEST TERMS. Compare in hours 21 hours to 3 days ..PLEASE HELP

OpenStudy (anonymous):

21:3 3 (number of days)*24 (hours in a day)=72 21:72 (number of hours in 3 days) 21/3=7 (reduced number of hours) 72/3= 24 (reduced number of hours) so the reduced ratio is 7:24

OpenStudy (anonymous):

Hope that helped :)